OpenStudy (anonymous):

A pair of equations is shown below: y = 7x − 8 y = 5x − 2 Part A: Explain how you will solve the pair of equations by substitution or elimination. Show all the steps and write the solution. Part B: If the two equations are graphed, at what point will the lines representing the two equations intersect? Explain your answer.

OpenStudy (princeharryyy):

7x - 8 = 5x -2 7x -5x-8 = 5x -5x -2 2x-8 = -2 2x-8+2 = -2+2 2x-6 = 0 x =3, y = 7*3 -8 = 13

OpenStudy (princeharryyy):

u can solve the second one by urself. I guess!

OpenStudy (anonymous):

ummmmmm

OpenStudy (anonymous):

can u guys just be nice?

OpenStudy (anonymous):

what? @SilverStream

OpenStudy (anonymous):

rlly u can help everyone with what they need help with not just cha cha cha

OpenStudy (anonymous):

^^? okay then, umm back to what you were saying, how do i know which is substitution or elimination? @princeharryyy

OpenStudy (princeharryyy):

try thinking you will get it by yourself else you can return back to me.

OpenStudy (anonymous):

i literally have no idea..

OpenStudy (princeharryyy):

ohk here it is literally.

OpenStudy (princeharryyy):

y = 7x-8 being put in equation 2 from equation one means substitution

OpenStudy (princeharryyy):

when we reduce the equation to find value what I did with x is elimination and when i found the value of x and again put it in equation 1 to find y that's sustitution

OpenStudy (princeharryyy):

substitution literally means to replace according to the dictionary definition. @paytoncobain I hope you got it.